OpenStudy (anonymous):

A tugboat goes 100 miles upstream in 10 hours. the return trip downstream takes 5 hours. find the speed of the tugboat without a current and the speed of the current

OpenStudy (anonymous):

distance = speed x time, so all you need to do is plug in the numbers for the distance and time of the first trip and solve for speed (by dividing distance by time). The same for the second trip.

OpenStudy (anonymous):

so the equation looks like 100=x(10)?

OpenStudy (anonymous):

Yes: 100 miles = speed * 10 hours \[\frac{ 100 miles }{ 10 hours } = speed\] 100/10 = 10, so 10 miles per hour.

OpenStudy (anonymous):

okay and then i got 20 for the second equation, did i do that correctly

OpenStudy (anonymous):

Yep.